The weight of teenagers in Barangay Payapa approaches a normal distribution with a mean of 52.9 kg and a standard deviation of 7.2 kg. If a teenager is a chosen at random from Barangay Payapa, what is the probability that his/her weight is above 48.5 kg?


1
Expert's answer
2022-04-13T18:07:23-0400

"P(X>48.5)=P(Z>\frac{48.5-52.9}{7.2})=P(Z>-0.61)=1-P(Z<-0.61)=0.7291."